Solve. s^2+3s-54+0
I think I use the quadratic equation to solve this. I did and got the answers of -3Plus or minus the square root of -207 all divided by 2. Is this right? Thank you for your help.

Answer by stanbon(75887)

s^2+3s-54 = 0
(s+9)(s-6) = 0
s = -9 or s = 6
=====================
Cheers,
Stan H.